A Career in Physics?

Of Course!

Pekka K. Sinervo Department of Physics

University of Toronto

1. What Is a Physicist?

2. Education, Training

3. A Few Typical Days...

4. Compensation and Career Development 5. Supply and Demand

6. How to Learn More

Upper Canada College February 1998

What is a Physicist?

The definition of a physicist is fuzzy – a few choices:

• Dictionary:

A specialist in physics, a science that deals with matter and energy and their interactions in the fields of mechanics, acoustics, optics, heat, electricity, magnetism, and radiation.

• Someone with a Ph.D., M.Sc. or B.Sc. in Physics, Physics and Mathematics, etc.

• Someone trained to observe/ study physical “phenomena”

– Phenomena related to the underlying structure of our world

– Separate from specialized areas such as Chemistry, Geology, or Astronomy

• Someone with the mathematical skills to interpret observations and place them in the context of a “model” or “theory”

Perhaps the most accurate definition is someone who

• wants to know how the world works

• use observational and theory to improve that understanding

⇒ A problem solver!

Problem Solving as a Career!

Physicists are basically technically literate problem solvers

• Have an advanced technical training

– Expected to know how to tackle pretty much any sort of new problem

– Have the know-how and experience to not be afraid of any tools

∗ Instrumentation for measurement or observation ∗ Electronics, lasers, particle detectors, sensors

∗ Data acquisition and analysis tools, computers

• Have training to think creatively and abstractly – Creativity is essential – way to generate “knowledge”

– Abstract thinking needed for “model-building” process – Observations are only useful to a physicist if they can be

used to fine-tune an existing model, or develop a new one

This will come clearer in a few examples

Jobs for Physicists

Physicists are found in a large variety of roles

• Traditional roles are in academia and research – Professors in universities ⇒ research & teaching

– Research scientists at universities or government labs

• More frequently found in the private sector

– Research scientists involved in

∗ Basic research focussed on certain technologies, eg. · HP laser printer technology · Tomography for MRI and PET scanners · Lasers for grocery scanners and CD players

∗ Technologies where principles “understood,” eg. · Sensors for oil well “logging” (Schlumberger-Doll) · Long range weather forecasting (Environment Canada) · Power transmission technology (Ontario Hydro)

∗ Modelling and analysis of “phenomena,” eg. · Nuclear reactor safety issues · Searching for mineral and oil deposits · Financial markets (Wall Street and Bay Street)!

– Probably takes in about 60-70% of trained physicists

Diversity of Physics

Physics is a very diverse field, eg. at U of T

• Quantum Optics (lasers!) – Making higher power lasers

– Making more precise lasers

• Condensed Matter Physics – properties of materials

– unique behaviour (eg. high Tc)

• Planetary Physics – Study of seismic waves

– Effect of glaciers on the earth’s surface

• Atmospheric Physics – creation of storms

– atmospheric pollution

– Effect of glaciers on the earth’s surface

• Subatomic Physics – exploring the interior of atoms

– Figuring out how our universe works

Laser Physicist

Lasers now appear to be ubiquitous

• Probably have at least one in your home or workplace... – CD players

– Some smoke detectors or alarm systems

– Grocery scanners

– Laser light shows....

• At U of T, have a large laser (“quantum optics”) group

– Do research in basic physical properties involved in lasing or using lasers

∗ Developing techniques for “trapping” atoms

∗ This line of research just won Nobel Prize in Physics last year

– Also working with various companies ∗ Lumonics Corporation has a “chair” in physics

∗ Working on developing high power lasers on the cheap

Department has five laser labs, each with many different set-ups looking at different phenomena

• Each optics table has different experiment going on

• May take months to get it right

Typical Day for a Quantum Optician...

Typical day involves

• Come in and warm up the research laser

– Look at last nights results – figure out what might have gone wrong

– Talk to supervisor or students

• Set up and do another set of measurements

– Try to understand results ∗ If they look OK, move on to next possible problem ∗ If not, think about what might still be wrong

– If they look “wierd,” think about how to confirm them ∗ Try to reproduce them

∗ Think about ways of mimicing the result

• Later in day, check to see if anything new has come into the library

– Have a break to talk to colleagues

– Check on results of other experiments

Typically are working on several things simultaneously

Another Typical Day – Particle Physicist

Come in and read e-mail

• I get typically about 30-40/day associated with research

• Catch up on stuff I can do immediately

Have weekly group meeting

• Meet and discuss progress in last week

• Bring up any issues or problems, eg.

– Out of disk space on local workstation cluster

– Somebody found software bug - has anybody else seen it?

– Schedule time to review a paper

• Go around the “table,” report on projects – Students and staff will report on their own research work – Keep everybody up-to-date

– Give everybody a chance to comment and advise

More on My Typical Day...

Spend couple hours in instrumentation lab

• Working on developing a small “inchworm” to very precisely precision a big detector

– Have had problems making it strong enough

– Worked with engineers and machinists to improve it • Also have several other “hardware” projects underway

– Work on microelectronic circuits

– New types of particle detectors

Later in day, meet 1-on-1 with other researchers to discuss progress

• Sometimes will sit down and go over data analysis

• Go over drafts of papers

Why So Appealing?

Always learning new things

• Advancing our knowledge about how world works

• Latest advances in high speed computing and communications

• New technologies

Interact with some very bright and pleasant people

• Not just a bunch of “nerds”

• Real people with similar passions about discovery

• Intellectually challenging

Often pretty competitive

• Want to be “first” to get there

• Want to be able to make progress the fastest

• Sometimes this is not so much fun – stressful!

Compensation and Career Development

Starting salaries for Ph.D. physicists vary great deal

• Research associates start around $30,000 to $45,000

• Assistant Professors start around $50,000 to $65,000 – Increase by about 2-6%/year depending on performance

• In industry, starting salaries are similar, but peak earlier

– Most large private employers recognize that research scientists should follow pattern similar to academia

Cost of getting a degree outweighed by higher salary

• But it does require some patience and devotion

• Can take up to 10 years of university education

Longer term career prospects are varied

• Academic stream is very stable on average

• Research scientists also have very low turn-over

• For B.Sc.’s in Physics, experience will vary depending on particular field

– Degree considered very useful

– Survey asking to rank usefulness between 1-5 ∗ Ph.D. on average ranked it as 4.2 ± 1.0

∗ M.Sc. on average ranked it as 4.0 ± 1.4

∗ B.Sc. on average ranked it as 3.1 ± 1.3

Education and Training

First step is OAC credits in MAT (all 3!) and PHY

• Must at least be comfortable with mathematics

• Curiousity and fooling around with gadgets is a plus

Second step is completion of B.Sc. in Physics, Math & Physics, Astronomy & Physics, etc.

• Provides experience and basic tools to function as a researcher or problem solver

• However, modest amount of real research experience

– People who complete at this stage go into technical areas

– Typically, are not involved in research-oriented careers

• Good training ground for people who want to become – technician in a research laboratory,

– enter a related postgraduate degree,

– a science or math teacher (note: ETP),

– involved in IT industry

Beyond the B.Sc.

Half of U of T students go on for postgraduate degrees

• Some complete 1 or 2-year M.Sc.

– Provides greater degree of hands-on research experience – Most students do not choose this path unless very spe-

cific goals

• Majority of our students go on for Ph.D. – Takes about 4-5 years

– Supported at the level of about $15,000-$22,000 in scholarships and research assistanceships

– Takes a lot of hard work...

∗ One to two years of course work

∗ Three to four years of research work

∗ Culminates in Ph.D. dissertation

Supply and Demand

Demand is important criteria in choosing a field

• Canada annually produces – about 650 B.Sc. physicists

– about 220 B.Sc. physicists

– about 120 Ph.D. physicists

• Unemployment rate is virtually unmeasurable – Appears to be less than about 2%

– Results come from recent surveys by CAP and NSERC • All of these indicate strong demand for people with Physics

degrees

This reflects the utility of a degree in Physics

• Professional problem solvers

• Trained to tackle toughest research problems

• Very versatile technically

Summary

Physics is not just a career – it is a way of thinking

• Allows one to be both creative and analytical

• Very stimulating and intellectually challenging

• Opens up a variety of doors in many fields

Remains the place where discoveries will continue to be made

• Academic research is still very challenging

• Private sector continues to play a significant role

Job prospects look quite good

• Strong demand both in academia and in industry

• Versatility is a big bonus in the private sector

• Employment rates are very high

Physics is both Fun and Challenging